OpenStudy (love_to_love_you):

Please help me! Will fan and give medal. Evaluate the expression for the given value of the variable. |-4b-8| + |-1-b^2| + 2b^3; b=-2

OpenStudy (kirbykirby):

\( |-4(-2)-8| + |-1-(-2)^2| + 2(-2)^3\) \( = |8-8|+|-1-(4)|+2(-8) \) can you figure the rest?

OpenStudy (love_to_love_you):

No, that's why I am asking for help. I have a really bad teacher :/

OpenStudy (kirbykirby):

Ok Well, just to be sure.. is there anything you didn't understand in the 2nd line I wrote? What I simplified was: \(-4(-2)=8\) \((-2)^2=4\) \((-2)^3=-8\)

OpenStudy (love_to_love_you):

No i understood it

OpenStudy (kirbykirby):

perfect, the next steps should be fairly simple then. \( = |8-8|+|-1-(4)|+2(-8)\) \( =|0|+|-1-4|-16\\ =0+|-5|-16\\ =5-16\\ =-11\) This is because: any number between absolute value bars | | will be a positive number. For example: |-5| = 5, and |5| = 5 |-2| = 2, etc...